# On Computational Limits and Provably Efficient Criteria of Visual Autoregressive Models: A Fine-Grained Complexity Analysis

### v1

@@ C.1 Lipschitz of Polynomial
Lemma C.1 (Lipschitz of polynomial).
@@
  • Let $M>1$.
  • Suppose we have $|x|\leq M,|x^{\prime}|\leq M$.
@@
$|f(x)-f(x^{\prime})|\leq O(M^{g-1})\cdot|x-x^{\prime}|$ (2)
@@
$\leq$ $~{}\sum_{i=1}^{g}|a_{i}\cdot(x-x^{\prime})|\cdot i\cdot M^{i-1}$
$=$ $~{}|x-x^{\prime}|\cdot\sum_{i=1}^{g}|a_{i}|\cdot i\cdot M^{i-1}$
$\leq$ $~{}O(M^{g-1})\cdot|x-x^{\prime}|$
The first step above is a consequence of the condition that $|x|\leq M$ and $|x^{\prime}|\leq M$, And the second derives from basic algebra, and the final step is due to $M>1$. ∎
C.2 Error Propagation of Inner Product
@@
Lemma C.2 (Error propagation of inner product).
@@
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
@@
$|\langle u^{\prime},v^{\prime}\rangle-\langle u,v\rangle|\leq 2k\epsilon M$
@@
$\leq$ $~{}\sum_{i=1}^{k}2\cdot M\cdot\epsilon$
$\leq$ $~{}2k\epsilon M$
The first step is derived from the conditions that each entry is at most $M$, $\|u^{\prime}-u\|_{\infty}\leq\epsilon$ and $\|v^{\prime}-v\|_{\infty}\leq\epsilon$. The second step follows directly from algebraic manipulation.
C.3 Error Analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{AAttC}(X)$
@@
Lemma C.3 (Error analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{AAttC}(X)$, formal version of Lemma 5.4).
@@
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
  • Let the polynomial approximation of attention matrix $\mathsf{AAttC}(X)$ as given in Definition 4.2.
@@
$\|\mathsf{AAttC}(X^{\prime})-\mathsf{AAttC}(X)\|_{\infty}\leq O(kM^{g+2}d)\cdot\epsilon$
@@
$\leq$ $~{}M\cdot d\cdot\epsilon$ (3)
The initial step is derived from the computation of matrix $Q$. The second step is a consequence of basic algebra, the third step arises from standard matrix multiplication, and the final step is a result of the condition $|X-X^{\prime}|_{\infty}\leq\epsilon$ and the fact that each entry is bounded by $M$.
In the same way, we can have $\|K-K^{\prime}\|_{\infty}\leq d\cdot\epsilon\cdot M$.
@@
$\|U-U^{\prime}\|_{\infty}\leq$ $~{}O(M^{g-1})\cdot\|Q-Q^{\prime}\|_{\infty}$
$\leq$ $~{}O(M^{g}d)\cdot\epsilon$
The first step above is derived from Lemma C.1 and the condition that each entry is bounded by $M$, while the second step results from Eq. (C.3) and basic algebra.
In the same way, we can have $\|V-V^{\prime}\|_{\infty}\leq O(M^{g}d)\cdot\epsilon$.
@@
$\leq$ $~{}2k\cdot O(M^{g}d)\cdot\epsilon\cdot M$
$\leq$ $~{}O(kM^{g+1}d)\cdot\epsilon$
The first step above is a result of basic algebra, the second step comes from Lemma C.2 and the lemma’s condition, and the final step is derived from basic algebra.
@@
$\|\mathsf{AAttC}(X^{\prime})-\mathsf{AAttC}(X)\|_{\infty}\leq O(kM^{g+1}d)\cdot\epsilon$
@@
C.4 Error Analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{Attn}(X)$
@@
Lemma C.4 (Error analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{Attn}(X)$, formal version of Lemma 5.5).
@@
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
  • Let the attention matrix $\mathsf{Attn}(X)\in\mathbb{R}^{n\times n}$ as given in Definition 3.3.
  • Let the polynomial approximation of attention matrix $\mathsf{AAttC}(X)\in\mathbb{R}^{n\times n}$ as given in Definition 4.2.
@@
$\|\mathsf{AAttC}(X^{\prime})-\mathsf{Attn}(X)\|_{\infty}\leq O(kM^{g+1}d)\cdot\epsilon$
@@
$\leq$ $~{}O(kM^{g+1}d)\cdot\epsilon+\epsilon$
$=$ $~{}O(kM^{g+1}d)\cdot\epsilon$
The first step is based on simple algebra, the second step is derived using the triangle inequality, the third step is obtained from Lemma C.3 and Lemma E.2, and the final step results from basic algebra.
C.5 Error Analysis of Up-Interpolation Layer
@@
Lemma C.5 (Error Analysis of Up-Interpolation Layer).
Assuming the conditions below are satisfied:
@@
  • Define $X^{\prime}\in\mathbb{R}^{h^{\prime}\times w^{\prime}\times d}$ as the approximated input feature map.
  • Let $\phi_{\rm up}:\mathbb{R}^{h\times w\times d}\to\mathbb{R}^{h^{\prime}\times w^{\prime}\times d}$ represent the up-interpolation layer defined in Definition 3.2.
@@
$\|\phi_{\rm up}(X)-\phi_{\rm up}(X^{\prime})\|_{\infty}\leq 16\epsilon$
@@
$\leq$ $~{}16\epsilon$
The first step is based on Definition 3.2, the second step is derived using the triangle inequality, the third step is a consequence of $|X-X^{\prime}|_{\infty}\leq\epsilon$, and the final step follows from $W(x)\in[0,1]$ and basic algebra.
@@
$\|\phi_{\rm up}(X)-\phi_{\rm up}(X^{\prime})\|_{\infty}\leq 16\epsilon$
Recall we have defined $\phi_{\mathrm{up}}:\mathbb{R}^{h\times w\times c}\to\mathbb{R}^{h^{\prime}\times w^{\prime}\times c}$ in Definition 3.2. Since there is no non-linear operation in $\phi_{\mathrm{up}}$, $\phi_{\mathrm{up}}$ is equivalent to a matrix multiplication operation, where the dimension of the matrix is $\mathbb{R}^{h^{\prime}w^{\prime}\times hw}$. For simplicity, we view $\phi_{\mathrm{up}}$ as a $\mathbb{R}^{h^{\prime}w^{\prime}\times hw}$ dimension matrix in the following proofs.
Remark C.6 (Applying $\phi_{\mathrm{up}}$ on $X\in\mathbb{R}^{n\times d}$).
The actual input of $\operatorname{\mathsf{VAR}}$ Transformer Layer are $r$ input token maps, $X_{1}\in\mathbb{R}^{h_{1}\times w_{1}\times d},\ldots,X_{r}\in\mathbb{R}^{h_{r}\times w_{r}\times d}$. We denote them as $X\in\mathbb{R}^{n\times d}$, where $n:=\sum_{i=1}^{r}h_{i}w_{i}$. We denote $\phi_{\mathrm{up}}(X)\in\mathbb{R}^{n^{\prime}\times d}$ as applying $\phi_{\mathrm{up}}$ to each $X_{i}\in\mathbb{R}^{h_{i}\times w_{i}\times d}$ for $i\in[r]$, where $n^{\prime}=\sum_{i=1}^{r}h_{i}^{\prime}w_{i}^{\prime}$.
C.6 Error Analysis for $\operatorname{\mathsf{VAR}}$ Transformer
@@
Lemma C.7 (Error propagation analysis for one $\operatorname{\mathsf{VAR}}$ Transformer Layer, formal version of Lemma 5.6).
@@
  • Let the $\operatorname{\mathsf{VAR}}$ Transformer Layer $\mathsf{F}_{\mathrm{var}}$ as given in Definition 3.5.
  • Let the Fast $\operatorname{\mathsf{VAR}}$ Transformer Layer $\mathsf{F}_{\rm fvar}$ as given in Definition 3.6.
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
  • Let the attention matrix $\mathsf{Attn}(X)\in\mathbb{R}^{n\times n}$ be defined as Definition 3.3.
  • Let the polynomial approximation of attention matrix $\mathsf{AAttC}(X)\in\mathbb{R}^{n\times n}$ as given in Definition 4.2.
@@
$\|\mathsf{F}_{\rm fvar}(X^{\prime})-\mathsf{F}_{\mathrm{var}}(X)\|_{\infty}\leq O(kM^{g+1}d)\cdot\epsilon$
@@
Firstly, we demonstrate that
$\|\mathsf{F}_{\rm fvar}(X^{\prime})-\mathsf{F}_{\mathrm{var}}(X)\|_{\infty}=$ $~{}\|\mathsf{AAttC}(\phi_{\rm up}(X^{\prime}))-\mathsf{Attn}(\phi_{\rm up}(X))\|_{\infty}$
$\leq$ $~{}O(kM^{g+1}d)\cdot 16\epsilon$
$\leq$ $~{}O(kM^{g+1}d)\cdot\epsilon$
The first step is derived from Definition 3.7 and Definition 3.6, the second step is obtained using Lemma C.5 and Lemma C.4, and the final step follows from basic algebra. ∎
C.7 Error Analysis for the Fast $\operatorname{\mathsf{VAR}}$ Transformer
@@
Lemma C.8 (Error analysis of the Fast $\operatorname{\mathsf{VAR}}$ Transformer, formal version of Lemma 5.7).
Assuming the conditions below are satisfied:
  • Let $X_{0}\in\mathbb{R}^{1\times d}$ denote the initial input data matrix.
  • Let the $\operatorname{\mathsf{VAR}}$ Transformer Layer $\mathsf{F}_{\mathrm{var}}$ be defined as Definition 3.5.
  • Let the Fast $\operatorname{\mathsf{VAR}}$ Transformer Layer $\mathsf{F}_{\mathrm{fvar}}$ as given in Definition 3.6.
  • Let $r\in\mathcal{N}$ represent the total iteration of the Fast $\operatorname{\mathsf{VAR}}$ model.
  • For $i\in[r]$, let $T_{i}(X)$ denote the intermediate variable of the $\operatorname{\mathsf{VAR}}$ model, as defined in Definition 3.7.
  • For $i\in[r]$, let $\widetilde{T}_{i}(X)$ denote the intermediate variable of the fast $\operatorname{\mathsf{VAR}}$ model, as defined in Definition 3.8.
  • Define $T_{r}(X)\in\mathbb{R}^{O(n^{2})\times d}$ a the final output of the $\operatorname{\mathsf{VAR}}$ Transformer.
  • Define $\widetilde{T}_{r}(X)\in\mathbb{R}^{O(n^{2})\times d}$ as the final output of the Fast $\operatorname{\mathsf{VAR}}$ Transformer.
@@
Then, we can show that the error bound of the final output $\widetilde{T}_{r}(X)$ as
$\|\widetilde{T}_{r}(X_{0})-T_{r}(X_{0})\|_{\infty}\leq 1/\operatorname{poly}(n)$
@@
$\|\widetilde{T}_{1}(X_{0})-T_{1}(X_{0})\|_{\infty}=$ $~{}\|\mathsf{F}_{\mathrm{fvar}}(X_{0})-\mathsf{F}_{\mathrm{var}}(X_{0})\|_{\infty}$
$\leq$ $~{}\|\mathsf{AAttC}(\phi_{\rm up}(X_{0}))-\mathsf{Attn}(\phi_{\rm up}(X_{0}))\|_{\infty}$
@@
The first step is based on Definition 3.8 and Definition 3.7, the second step is derived from Definition 3.6 and Definition 3.5, and the final step follows from Lemma E.2.
Assume that the following statement is true for the $k$-th iteration:
$\|\widetilde{T}_{k}(X_{0})-T_{k}(X_{0})\|_{\infty}\leq 1/\operatorname{poly}(n)$ (4)
@@
$\|\widetilde{T}_{k+1}(X_{0})-T_{k+1}(X_{0})\|_{\infty}=$ $~{}\|\mathsf{F}_{\mathrm{fvar}}(\widetilde{T}_{k}(X_{0}))-\mathsf{F}_{\mathrm{var}}(T_{k}(X_{0}))\|_{\infty}$
$\leq$ $~{}O(k\cdot(\operatorname{poly}(n))^{g+1}\cdot d)\cdot(1/\operatorname{poly}(n))$
@@
The first step is based on Definition 3.8 and Definition 3.7, the second step is derived from Lemma C.7, the fact that each entry in the matrices can be represented using $O(\log(n))$ bits, and Eq. (4), while the final step results from basic algebra.
@@
$\|\widetilde{T}_{r}(X_{0})-T_{r}(X_{0})\|_{\infty}\leq 1/\operatorname{poly}(n).$
@@
Appendix D Error Analysis of VQVAE Decoder

### v2

@@ B.1 Lipschitz of Polynomial
Lemma B.1 (Lipschitz of polynomial).
@@
  • Let $R>1$.
  • Suppose we have $|x|\leq R,|x^{\prime}|\leq R$.
@@
$|f(x)-f(x^{\prime})|\leq O(R^{g-1})\cdot|x-x^{\prime}|$ (2)
@@
$\leq$ $~{}\sum_{i=1}^{g}|a_{i}\cdot(x-x^{\prime})|\cdot i\cdot R^{i-1}$
$=$ $~{}|x-x^{\prime}|\cdot\sum_{i=1}^{g}|a_{i}|\cdot i\cdot R^{i-1}$
$\leq$ $~{}O(R^{g-1})\cdot|x-x^{\prime}|$
The first step above is a consequence of the condition that $|x|\leq R$ and $|x^{\prime}|\leq R$, And the second and last step derive from basic algebra. ∎
B.2 Error Propagation of Inner Product
@@
Lemma B.2 (Error propagation of inner product).
@@
  • Let $R>1$.
  • Assume the value of each entry in matrices can be bounded by $R$.
@@
$|\langle u^{\prime},v^{\prime}\rangle-\langle u,v\rangle|\leq 2k\epsilon R$
@@
$\leq$ $~{}\sum_{i=1}^{k}2\cdot R\cdot\epsilon$
$\leq$ $~{}2k\epsilon R$
The first step is derived from the conditions that each entry is at most $R$, $\|u^{\prime}-u\|_{\infty}\leq\epsilon$ and $\|v^{\prime}-v\|_{\infty}\leq\epsilon$. The second step follows directly from algebraic manipulation.
B.3 Error Analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{AAttC}(X)$
@@
Lemma B.3 (Error analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{AAttC}(X)$).
@@
  • Let $R>1$.
  • Assume the value of each entry in matrices can be bounded by $R$.
  • Let $\mathsf{AAttC}$ denote the approximated attention layer defined in Definition 4.2.
@@
$\|\mathsf{AAttC}(X^{\prime})-\mathsf{AAttC}(X)\|_{\infty}\leq O(kR^{g+2}d)\cdot\epsilon$
@@
$\leq$ $~{}R\cdot d\cdot\epsilon$ (3)
The initial step is derived from the computation of matrix $Q$. The second step is a consequence of basic algebra, the third step arises from standard matrix multiplication, and the final step is a result of the condition $|X-X^{\prime}|_{\infty}\leq\epsilon$ and the fact that each entry is bounded by $R$.
In the same way, we can have $\|K-K^{\prime}\|_{\infty}\leq d\cdot\epsilon\cdot R$.
@@
$\|U-U^{\prime}\|_{\infty}\leq$ $~{}O(R^{g-1})\cdot\|Q-Q^{\prime}\|_{\infty}$
$\leq$ $~{}O(R^{g}d)\cdot\epsilon$
The first step above is derived from Lemma B.1 and the condition that each entry is bounded by $R$, while the second step results from Eq. (B.3) and basic algebra.
In the same way, we can have $\|V-V^{\prime}\|_{\infty}\leq O(R^{g}d)\cdot\epsilon$.
@@
$\leq$ $~{}2k\cdot O(R^{g}d)\cdot\epsilon\cdot R$
$\leq$ $~{}O(kR^{g+1}d)\cdot\epsilon$
The first step above is a result of basic algebra, the second step comes from Lemma B.2 and the lemma’s condition, and the final step is derived from basic algebra.
@@
$\|\mathsf{AAttC}(X^{\prime})-\mathsf{AAttC}(X)\|_{\infty}\leq O(kR^{g+1}d)\cdot\epsilon$
@@
B.4 Error Analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{Attn}(X)$
@@
Lemma B.4 (Error analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{Attn}(X)$).
@@
  • Let $R>1$.
  • Assume the value of each entry in matrices can be bounded by $R$.
  • Let $\mathsf{Attn}$ denote the attention layer defined in Definition 3.4.
  • Let $\mathsf{AAttC}$ denote the approximated attention layer defined in Definition 4.2.
@@
$\|\mathsf{AAttC}(X^{\prime})-\mathsf{Attn}(X)\|_{\infty}\leq O(kR^{g+1}d)\cdot\epsilon$
@@
$\leq$ $~{}O(kR^{g+1}d)\cdot\epsilon+\epsilon$
$=$ $~{}O(kR^{g+1}d)\cdot\epsilon$
The first step is based on simple algebra, the second step is derived using the triangle inequality, the third step is obtained from Lemma B.3 and Lemma D.2, and the final step results from basic algebra.
B.5 Error Analysis of Up-Interpolation Layer
@@
Lemma B.5 (Error Analysis of Up-Interpolation Layer).
If the following conditions hold:
@@
  • Define $X^{\prime}\in\mathbb{R}^{h\times w\times d}$ as the approximated input feature map.
  • Let $\Phi_{\rm up}:\mathbb{R}^{h\times w\times d}\to\mathbb{R}^{h^{\prime}\times w^{\prime}\times d}$ represent the pyramid up-interpolation layer defined in Definition 3.3.
@@
$\|\Phi_{\rm up}(X)-\Phi_{\rm up}(X^{\prime})\|_{\infty}\leq O(\epsilon)$
@@
$=$ $~{}O(\epsilon)$
The first step is based on Definition 3.2, the second step is derived using the triangle inequality, the third step is a consequence of $\|X-X^{\prime}\|_{\infty}\leq\epsilon$, and the final step follows from $W(x)\in[0,1]$ and basic algebra.
@@
$\|\Phi_{\rm up}(X)-\Phi_{\rm up}(X^{\prime})\|_{\infty}\leq O(\epsilon)$
B.6 Error Analysis for $\operatorname{\mathsf{VAR}}$ Transformer
@@
Lemma B.6 (Error propagation analysis for one $\operatorname{\mathsf{VAR}}$ Transformer Layer).
@@
  • Let $R>1$.
  • Assume the value of each entry in matrices can be bounded by $R$.
  • Let $\mathsf{Attn}$ denote the attention layer defined in Definition 3.4.
  • Let $\mathsf{AAttC}$ denote the approximated attention layer defined in Definition 4.2.
@@
$\|\mathsf{AAttC}(\Phi_{\rm up}(X^{\prime}))-\mathsf{Attn}(\Phi_{\rm up}(X))\|_{\infty}\leq O(kR^{g+1}d)\cdot\epsilon$
@@
The result is easily derived from Lemma B.4 and Lemma B.5. ∎
B.7 Error Analysis for the Fast $\operatorname{\mathsf{VAR}}$ Transformer
@@
Lemma B.7 (Error analysis of the Fast $\operatorname{\mathsf{VAR}}$ Transformer, formal version of Lemma 5.4).
If the following conditions hold:
  • Let $X_{\mathrm{init}}\in\mathbb{R}^{1\times d}$ denote the initial input token map.
  • Let $K\in\mathbb{N}$ represent the number of approximate attention layers in the $\operatorname{\mathsf{VAR}}$ model.
  • Let the $\operatorname{\mathsf{VAR}}$ transformer $\mathsf{TF}_{K}$ be defined as Definition 3.5.
  • Let the Fast $\operatorname{\mathsf{VAR}}$ Transformer $\mathsf{FTF}_{K}$ as given in Definition 4.3.
  • For $i\in[K]$, let $\mathsf{TF}_{i}(X_{\mathrm{init}})$ denote the output of the i-th iteration of the $\operatorname{\mathsf{VAR}}$ transformer.
  • For $i\in[K]$, let $\mathsf{FTF}_{i}(X_{\mathrm{init}})$ denote the output of the i-th iteration of the fast $\operatorname{\mathsf{VAR}}$ transformer.
  • Let $\mathsf{TF}_{K}(X_{\mathrm{init}})\in\mathbb{R}^{O(n^{2})\times d}$ denote the final output of the $\operatorname{\mathsf{VAR}}$ transformer.
  • Let $\mathsf{FTF}_{K}(X_{\mathrm{init}})\in\mathbb{R}^{O(n^{2})\times d}$ denote the final output of the fast $\operatorname{\mathsf{VAR}}$ transformer.
@@
Then, we can show that the error bound of the final output $\mathsf{FTF}_{K}(X_{\mathrm{init}})$ as
$\|\mathsf{FTF}_{K}(X_{\mathrm{init}})-\mathsf{TF}_{K}(X_{\mathrm{init}})\|_{\infty}\leq 1/\operatorname{poly}(n)$
@@
$\|\mathsf{FTF}_{1}(X_{\mathrm{init}})-\mathsf{TF}_{1}(X_{\mathrm{init}})\|_{\infty}=$ $~{}\|\mathsf{AAttC}_{1}(X_{\mathrm{init}})-\mathsf{Attn}_{1}(X_{\mathrm{init}})\|_{\infty}$
@@
The first step is based on Definition 3.5 and Definition 4.3, and the final step follows from Lemma D.2.
Assume that the following statement is true for the $k$-th iteration (where $k<K$):
$\|\mathsf{FTF}_{k}(X_{\mathrm{init}})-\mathsf{TF}_{k}(X_{\mathrm{init}})\|_{\infty}\leq 1/\operatorname{poly}(n)$ (4)
@@
$\|\mathsf{FTF}_{k+1}(X_{\mathrm{init}})-\mathsf{TF}_{k+1}(X_{\mathrm{init}})\|_{\infty}=$ $~{}\|\mathsf{AAttC}_{k+1}(\Phi_{\mathrm{up},k}(\mathsf{FTF}_{k}(X_{\mathrm{init}})))-\mathsf{Attn}_{k+1}(\Phi_{\mathrm{up},k}(\mathsf{TF}_{k}(X_{\mathrm{init}})))\|_{\infty}$
@@
The first step is based on Definition 3.5 and Definition 4.3, the second step is derived from Lemma B.6, the fact that each entry in the matrices can be represented using $O(\log(n))$ bits, and Eq. (4).
@@
$\|\mathsf{FTF}_{K}(X_{\mathrm{init}})-\mathsf{TF}_{K}(X_{\mathrm{init}})\|_{\infty}\leq 1/\operatorname{poly}(n)$
@@
Appendix C Error Analysis of VQVAE Decoder